Two thin, spherical, concentric metal shells have a radius of 6 cm and 8 cm. The inner shell carries a charge of 7 nC and the outer shell carries a charge of 4 nC. Find the resulting electric field at a) 7 cm from the centre; b) 10 cm from the centre.


1
Expert's answer
2022-03-03T12:05:30-0500

a)

"E_1=k(\frac{q_1(r-R_1)}{r_1^3}-\frac{q_2(R_2-r)}{r_2^3})=0.22~\frac Vm,"

b)

"E_2=\frac{kq_2R_2}{r_2^3}=0.7~\frac Vm ."